Solutions for y'= x²+y²/2xy
1
Expert's answer
2020-12-04T07:45:56-0500

Let us solve the differential equation "y'= \frac{x²+y²}{2xy}".


Let "y=ux". Then "y'=u'x+u". Consequently we have the following differential equation:


"u'x+u=\frac{x²+u²x^2}{2ux^2}=\frac{1+u²}{2u}"


"u'x=\frac{1+u²}{2u}-u=\frac{1-u²}{2u}"


"\frac{du}{dx}x=\frac{1-u²}{2u}"


"\frac{2udu}{1-u^2}=\frac{dx}{x}"


"\int\frac{2udu}{1-u^2}=\int\frac{dx}{x}"


"-\int\frac{d(1-u^2)}{1-u^2}=\int\frac{dx}{x}"


"-\ln|1-u^2|=\ln|x|-\ln|C|"


"\ln|C|=\ln|x(1-u^2)|"


"C=x(1-u^2)"


"C=x(1-\frac{y^2}{x^2})"


Therefore, the solution of the differential equation "y'= \frac{x²+y²}{2xy}" is the following:


"C=x-\frac{y^2}{x}"